Shift report 20011010 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Oct 10 2001 - 06:11:40 EDT

  • Next message: Flemming Videbaek: "Re: Shift report 20011010 00:00-08:00"

    ------------------------------------------------------------
    Date: Wed Oct 10 06:11:40 2001, Shift: 00:00-08:00
    Supervisor: Michael Murray, with: none
    Log book#9 pages 134 - 136
    
    Shift Summary:
    ------------------------------------------------------------
    1) There was beam from 21.30 or so last night until 1.40am
    The next ramp was aborted because of a quench and it took
    until 5.05 for the cryogenics to be ready. The second 
    attempt to ramp failed because of a "radiation monitor 
    failure/interlock". By this time the AGS had already shut
    down for the maintenance period and so Lee at MCR gave up. 
    
    2) Run 5276 was finished with 313k events.
    TPC Pedestal run 5277 and ordinary pedestal run 5278.
    
    3) There were no detector problems. Three red alarm lights
    in rack 8 were on but I did not know what they meant. 
    Towards the end of run 5276 D5 tripped because of high
    temperture. The magnet watch was able to reset it but 
    could not find the time of the trip because it is not 
    recorded well in the on screen error log of the magnet 
    control panel in building 1002B.
    Since there is no chance of beam for 2 days I turned down
    the magnets, set them to standby and went to bed.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Oct 10 2001 - 06:12:08 EDT